Two marbles are drawn without replacement from a box of 3 white, 2 green, and 1 blue marble. Find the probability that both marbles are white. Answer must be in fraction form.
Answer by stanbon(48568)

"Without replacement" is a key phrase here.
P(1st draw is white)= 3/6
P(2nd draw is white)=2/5
P(WW without replacement)= (3/6)(2/5)=1/5
Cheers,
Stan H.